OpenStudy (anonymous):

what is the slop of the line passing through the points ( 2, 7 ) and ( -1, 3 )

OpenStudy (anonymous):

Slope of line in two pint form is given as: \[m= \frac{y_2-y_1}{x_2-x_1} =\] \[\rightarrow m=\frac{3-7}{-1-2} =\frac{-4}{-3} =\frac{4}{3} \] is the required slope. @kagome